OpenStudy (anonymous):

A chemist needs to make 1000 mL of a 30% alcohol solution by mixing 25% and 55% solutions. How much of each should she use?

OpenStudy (mathstudent55):

Let the amount of 25% solution be x Let the amount of 55% solution be y To make 1000 ml, you have x + y = 1000 Now look at how much alcohol you have 0.25x + 0.55y = 0.3(1000) Solve these two equations simultaneously for x and y.

OpenStudy (anonymous):

would you plug in 1000-x for y in the second equation?

OpenStudy (anonymous):

nevermind i got it! thank you!